OpenStudy (anonymous):

3/2x-4=16 How do i solve for x ?

OpenStudy (ghazi):

multiply both sides by 2x-4

OpenStudy (ghazi):

\[(2x-4)*\frac{ 3 }{ 2x-4 }=16*(2x-4)\]

OpenStudy (ghazi):

now you'll have \[3=16(2x-4)=32x-64\]

OpenStudy (ghazi):

can you solve for x now?

OpenStudy (anonymous):

 thats what it was supposed to say sorry

OpenStudy (ghazi):

okay....multiply every term by 2x

OpenStudy (ghazi):

\[2x*\frac{ 3 }{ 2x }-4*2x=16*2x\]

OpenStudy (anonymous):

ok thanks

OpenStudy (ghazi):

:)

OpenStudy (ghazi):

3-8x=32x